The light beams of intensities in the ratio of 9 : 1 are allowed to interfere. What will be the ratio of the intensities of maxima and
minima ?
(A) 3 : 1 (B) 4 : 1 (C) 25 : 9 (D) 81 : 1

Grade:

2 Answers

Arun
3 years ago
Intensity of the wave is directly proportional to square of amplitude.
I ∝ a^2
I1/I2 = (a1/a2)^2
9/1 = (a1/a2)^2
a1/a2 = 3
a1 = 3a2
Ratio of maximum to minimum intensity by -
Imax / Imin = (a1+a2)^2 / (a1-a2)^2
Imax / Imin = (3a2+a2)^2 / (3a2-a2)^2
Imax / Imin = 4^2 / 2^2
Imax / Imin = 16 / 4
Imax / Imin = 4
Therefore, ratio of maximum to minimum intensity is 4:1.
Hope this helps you..
 
Vikas TU
3 years ago
The light beams with intensities being in the ratio 19, so the amplitudes are in the ratio 13. So, the maximum and minimum intensities are in the ratio (3−1)2(3+1)2=2242=416=14.